# Evaluation of Dataframe Libraries for Data Preparation on a Single Machine

### v1

@@ II Dataframes
III Bento
III-A Datasets
To evaluate the performance of the dataframe libraries on general, real-world data preparation use cases, we selected four datasets that cover a variety of sizes, complexities, and characteristics, ensuring a comprehensive evaluation of the libraries: (i) Athlete [46], collecting 0.2 million records covering the competition results achieved by the athletes through 120 years of Olympic Games history; (ii) Loan [47], containing 2 million rows about loan applicants and their financial profiles from the LendingClub company; (iii) Patrol [48], composed of 27 million records about 11 years of traffic stops by California law enforcement agencies; (iv) Taxi [49], containing 77 million records about taxi trips in New York City in 2015, hence also suitable for evaluating scalability. A summary of their main features is displayed in Table III.
Collected from Kaggle [25], all of these datasets were previously adopted in related work about data preparation and dataframe libraries. For instance, the three largest datasets are the ones selected by Petersohn et al. [21] to evaluate the performance of Modin, due to the variety of their features. Instead, Athlete was picked by Hameed and Naumann [4] to analyze how different preparators are supported by commercial data preparation tools.
@@ III-A Datasets
Figure 1: Average runtime for each stage of the data preparation pipeline (EDA, DT, DC) on different datasets.
III-B Data Preparation Pipelines
Overall, the pipelines employ a set of 27 preparators (reported in Table II) that can be clustered into four main stages:
  • Input/output (I/O): preparators that handle the input/output of data in various formats, such as databases, CSV or Parquet [50] files, Web APIs, etc.
  • Exploratory data analysis (EDA): preparators that support the exploratory analysis of the features of the data, for a better understanding of the data at hand and the detection of errors or anomalies.
  • Data transformation (DT): preparators that transform the data to make it more suitable for analysis, such as categorical encoding, join, aggregation, etc.
  • Data cleaning (DC): preparators to ensure the quality of the data at hand, thus the reliability of the results of its analysis (e.g., handling missing or erroneous data and correcting or removing outliers).
III-C Pipeline Specification
Bento is a Python framework that provides a general shared interface to design the data preparation pipeline, ignoring the differences in the APIs of the considered libraries. For preparator names, we adopted the convention proposed by Hameed and Naumann [4]. Each pipeline can be defined through a JSON file by specifying the sequence of preparators to employ. Our tool automatically deploys the pipeline for every specific library.
[…]
Unless differently stated, we load data from CSV files and measure the execution time under three key settings: (i) function-core, which evaluates each preparator alone (if the library adopts lazy evaluation, it requires to force the execution after each API call); (ii) pipeline-stage, which measures the runtime for the execution of each stage (i.e., I/O, EDA, DT, and DC); (iii) pipeline-full, which measures the runtime for the execution of the entire pipeline. Finally, our framework also supports the pipeline execution in Docker containers. Docker provides benefits like specifying core and memory usage and isolating Python environments to prevent package conflicts.
@@ III-C Pipeline Specification
Figure 2: Speedup over Pandas for each preparator on Athlete and Loan.
III-D Hardware and Software
All measurements are obtained as the average value over three runs on a machine equipped with an Intel Xeon Silver 4116 CPU (24 threads running at 2.1 GHz) and 196 GB of RAM, using Python 3.9, when not specified otherwise. The machine also features a NVIDIA Tesla T4 GPU with 16 GB of RAM and CUDA 11.2. The Ray engine was configured with default settings, resulting in 24 worker threads running. The Dask engine was configured comparably, leading to the creation of 6 workers and 24 threads for each execution. To ensure accuracy and avoid warm-up overhead, the assessment of execution time occurs once the system (e.g., JVM) has completed its warming-up process. To assess the performance of the libraries across machines with different hardware specifications, we simulated three distinct machine configurations using Docker, as detailed in Table IV. Finally, we evaluated scalability on incremental samples of Taxi and Patrol.
(a) Patrol
(b) Taxi
Figure 3: Speedup over Pandas for each preparator on Patrol and Taxi.
Figure 4: Average runtime for the apply preparator on Patrol and Taxi (only libraries not raising OoM issues).
IV Evaluation Results
In this section, we report and analyze the detailed results of our extensive experimental evaluation of the presented dataframe libraries. Our main goal is to provide data scientists and practitiones with useful insights for supporting them in the selection of the best solution for their data preparation tasks. Thus, our comparison is designed to assess the performance of the dataframe libraries based on the operations to carry out (considering both the distinct preparators and the benefits introduced by lazy evaluation, when supported), the size and the features of the dataset at hand, and the configuration of the machine on which the pipeline is executed.
In particular, we organize our evaluation in multiple subsections, each of them is designed to answer one of the following research questions:
  1. Q1.
    What is the performance of the dataframe libraries in running data preparation pipelines (composed of the preparators applied most frequently by the Kaggle data science community) on datasets of different size and features? (Section IV-A)
  2. Q2.
    How does the optimizations introduced by lazy evaluations impact on the performance of the libraries that support it? (Section IV-B)
  3. Q3.
    How do the libraries scale by increasing the size of the datasets and varying the configuration of the machine used to run the data preparation pipeline (i.e., from laptop to server)? (Section IV-C)
IV-A Evaluation on Data Preparation Pipelines
Summary—For EDA, Polars is generally the best performer, up to 10 times faster than Pandas. For DT and DC, if a GPU is available, CuDF generally outperforms other libraries, but SparkSQL and Vaex achieve the best results on the largest dataset. Despite not supporting the Parquet format, DataTable is the fastest library for I/O.
For each of the selected datasets (Athlete, Loan, Patrol, Taxi), Figure 1 shows the runtime for the execution of three main stages of the data preparation pipeline: EDA, DT, and DC. The I/O stage is considered separately in Figures 6 and 6. Figure 1 considers each stage in its entirety, allowing to perform lazy evaluation at stage level when supported. Additionally, in Figures 2 and 3, we present the independent performance of each distinct preparator (i.e., we force the execution separately for each preparator), normalized with respect to Pandas (the de facto standard library for managing dataframes). For each preparator, the red line represents therefore Pandas as the baseline, while distinct markers denote the speedup for each library, which is defined as follows:
[…]
Figure 5: Average runtime for reading CSV and Parquet files.
@@ IV-A Evaluation on Data Preparation Pipelines
(a) Athlete, Write
(b) Loan, Write
(c) Patrol, Write
(d) Taxi, Write
Figure 5: Average runtime for reading CSV and Parquet files.
Figure 6: Average runtime for writing CSV and Parquet files.
IV-A1 Explorative Data Analysis (EDA)
In this stage, Polars clearly stands out as the best performer on all datasets but Taxi, where ModinD outperforms other dataframe libraries (Figures 0(a)-0(d)). In this stage, four preparators alone account for 95% of the total execution time: locate missing values (isna), locate outliers (outlier), search by pattern (srchptn), and sort values (sort).
Note that Polars registers the best performance for two of them, locate missing values and locate outliers (Figures 2 and 3), where it is about ${10^{4}}$ times faster than Pandas and 0.3 times faster than the second best performer, DataTable. In fact, Polars can quickly locate missing values by exploiting the specific metadata stored by Arrow arrays [51]. Nevertheless, DataTable records a better performance in locating missing values on Patrol (Figure 2(a)), with a speedup of ${10^{5}}$ over Pandas and ${10^{3}}$ over Polars. As highlighted by Table III, Patrol is mostly composed of string attributes. For this type of data, Polars uses Arrow large_utf8 buffers, which are more expensive to filter than Python strings/chars, requiring therefore extra work.
For the other two most relevant preparators, in Figure 2(a) we can observe that Vaex excels in searching by pattern, being about 100 times faster than Pandas and 10 times faster than CuDF, which is the second best performer. Vaex tends to be very efficient in column-wise operations due to its internal columnar organization of the data. On the other hand, it is much less efficient when applying a row-wise preparator, for instance in locating missing values or outliers, where it is even about 100 times slower than Pandas. For sorting values, while CuDF is the fastest library on Athlete, PySpark (in both versions) shows better results as the datasets get larger. Note that Modin needs to convert the dataframe back and forth to Pandas for the sorting operation, hence it is generally the worst performer, up to 100 times slower than SparkSQL.
Finally, including preparators accounting for less than 5% of the total execution time, such as get columns list (gcols), get columns types (dtypes), get dataframe statistics (stats), and query columns (query), Modin outperforms other libraries in computing statistics exploiting its multithreading power, which even makes ModinD faster than Polars when running the entire stage on Taxi. Overall, Pandas maintains good performance throughout the stage, while Pandas2 shows slight but not significant improvements over its predecessor.
IV-A2 Data Trasformation (DT)
In this stage, CuDF emerges as the best performer on Athlete (Figure 0(e)) and Patrol (Figure 0(g)), while Polars and SparkSQL register the best performance (also thanks to lazy evaluation) on Loan (Figure 0(f)) and Taxi (Figure 0(h)), respectively.
All libraries show better memory management than Pandas, which goes out of memory even on Patrol, with the calculate column using expressions (apply) preparator. The comparison among the libraries that successfully complete their execution is shown in Figure 4, considering their absolute runtime (i.e., the lower the better). Except for Vaex, all libraries struggle with column-wise operations, where Vaex outperforms them by leveraging its internal columnar organization. Note that when running the entire stage (Figure 0(g)) Pandas has enough time to release memory, preventing out-of-memory issues. Pandas2 follows the trend of its predecessor, without any significant improvement.
Finally, the lazy execution of SparkSQL proves to be extremely efficient when running the entire data transformation stage on larger datasets. In fact, it almost reaches the performance of CuDF on Patrol (Figure 0(g)), with a speedup of ${10^{3}}$ over Pandas, and emerges as the best performer on Taxi (Figure 0(h)), where Pandas (along with Pandas2 and SparkPD), Modin, and Polars run into out-of-memory issues.
IV-A3 Data Cleaning (DC)
@@
In this stage, the libraries that stand out the most are CuDF, and Vaex. In particular, CuDF consistently outperforms other libraries on Athlete and Loan, the two smallest datasets (Figures 0(i) and 0(j)). The pipelines designed for these two datasets include a preparator, deduplicate rows (dedup), accounting for 70% of the overall execution time for the data cleaning stage. Note that on Athlete CuDF is the only library performing better than Pandas (approximately 10 times faster), which on average maintains good performance, both in its basic version and as Pandas2. On Loan, it is possible to notice how Vaex equals or (even significantly) exceeds the performance of CuDF for multiple preparators, but the greater overall impact of the ones for which it is outperformed by CuDF, i.e., deduplicate rows, fill empty cells (fillna), and change date & time format (chdate), lead to a consistent difference in the overall execution time for this stage (Figure 0(j)).
@@
However, results are different when moving to larger datasets (Figures 0(k) and 0(l)). CuDF excels on Patrol, where it is over 100 times faster than Pandas in both preparators, i.e., delete empty rows (dropna) and change date & time format (chdate), and slightly outperforming Vaex (in dropna) and CuDF. Nevertheless, its performance decrease on Taxi, where instead Vaex achieves the best results, being 10 times faster than it and more than 100 times faster than Pandas.
@@
Figure 7: Average runtime for the execution of the entire data preparation pipelines.
IV-A4 Input/Output (I/O)
Loading data represents a fundamental step for the success of any data-driven process. In particular, CSV has established itself as the de facto standard for reading and writing raw data, and Pandas is recognized as an efficient and effective tool for its ingestion [52]. Nowadays, the diffusion of Arrow (whose adoption represents the main novelty of Pandas2 compared to its predecessor, as stated in Section II) is pushing towards an increasing adoption of the Parquet format [50], known for its efficient data compression. In particular, developers often load Parquet data into memory and convert it to Arrow [53], recognized as an ideal in-memory transport layer for such data [54].
Figures 6 and 6 show performance in read and write operations on both file formats, CSV and Parquet. DataTable, which maps the file into memory and navigates through it using pointers, records the best overall performance in reading files, but it does not support the Parquet format. Polars is the best performer in reading Parquet files due to its capability to read them directly into memory using Arrow. When dealing with small datasets like Athlete, Parquet and CSV exhibit almost the same performance (Figure 4(a)); in some cases (e.g., Vaex), the use of Parquet can even lead to worse performance, due to the need to decompress the data. When the size increases, the use of Parquet generally produces better results; for instance, reading Parquet files with SparkSQL is over 10 times faster than reading CSV files (Figure 4(d)).
Write operations exhibit a wide range of performance, mainly due to the need of compressing data before saving. Parquet files come with additional metadata and encoding, which may introduce some overhead for the writing process. Parquet performs better than CSV for smaller datasets, but this advantage often diminishes as the dataset size increases. For instance, while on Patrol the Parquet format brings significant advantages to Pandas (outperforming CSV by a factor of 27) thanks to its efficient dataframe serialization, Modin and CuDF are much faster in writing CSV files (Figure 5(c)). Nevertheless, CSV format causes CuDF to run out of memory on the largest dataset (Figure 5(d)), while it succeeds in writing Parquet files.
TABLE IV: Specifications of each machine configuration.
@@ IV-A4 Input/Output (I/O)
IV-B Impact of Lazy Evaluation
SummaryCuDF significantly outperforms other libraries when running the entire pipeline, but SparkSQL excels on the largest dataset. Lazy evaluation brings benefits mostly on large datasets, improving performance by 70% on average (up to almost 130% in some cases) compared to its eager counterpart.
Figure 7 reports for each library the runtime for the execution of the entire data preparation pipeline. For libraries supporting lazy evaluation (i.e., SparkPD, SparkSQL and Polars), we also point out the difference compared to eager evaluation. Considering the overall performance on the entire pipeline, CuDF consistently excels, by effectively leveraging on the power of GPU. Pandas performs well on smaller datasets, but struggles on larger ones, running out of memory even on Patrol. This is due to its eager approach (which requires to calculate intermediate results), causing higher memory consumption [55]. Additionally, Vaex and DataTable perform well on Patrol thanks to parallel processing, which allows them to efficiently exploit available CPU cores and manage many string columns.
Lazy evaluation brings significant benefits on large datasets, such as Patrol (Figure 6(c)). SparkPD takes advantage of this strategy with Pandas-specific optimizations (e.g., vectorized operations and memory efficiency), improving performance by 62%. Further, the good performance of Polars can be attributed to its query optimization through streaming processing, early filters, and projection pushdown, resulting in a remarkable 126% performance improvement over its eager counterpart (Figure 6(c)). Finally, it can be noticed how long logical plans can limit the optimization introduced by lazy evaluation. Yet, the execution plan overhead of SparkSQL does not yield substantial improvements for either smaller or larger datasets, with an improvement on Patrol limited to 31% and performance on Taxi almost unchanged.

### v2

@@ 3. Bento
Data Preparation Pipelines
Overall, the pipelines employ a set of 27 preparators (reported in Table 2) that can be clustered into four main stages:
  • Input/output (I/O): preparators to handle the input/output of data in various formats, such as databases, CSV or Parquet files, Web APIs, etc.
  • Exploratory data analysis (EDA): preparators for supporting the exploratory analysis of data features, to better understand the data at hand and detect errors or anomalies.
  • Data transformation (DT): preparators for transforming the data to make it more suitable for analysis, such as categorical encoding, join, aggregation, etc.
  • Data cleaning (DC): preparators to improve the quality of the data, hence the reliability of the results of its analysis (e.g., handling missing or erroneous data and correcting or removing outliers).
Pipeline Specification
Bento is a Python framework that provides a general shared interface to design the data preparation pipeline, ignoring the differences in the APIs of the considered libraries. For preparator names, we adopted the convention proposed by Hameed and Naumann (Hameed and Naumann, 2020). Each pipeline can be defined through a JSON file by specifying the sequence of preparators to employ. Our tool automatically deploys the pipeline for every specific library.
[…]
Unless differently stated, we load data from CSV files and measure the execution time under three key settings: (i) function-core, which evaluates each preparator alone (if the library adopts lazy evaluation, it requires to force the execution after each API call); (ii) pipeline-stage, which measures the runtime for the execution of each stage (i.e., I/O, EDA, DT, and DC); (iii) pipeline-full, which measures the runtime for the execution of the entire pipeline. Finally, our framework also supports the pipeline execution in Docker containers, which allow to specify core and memory usage and isolate Python environments to prevent package conflicts.
Hardware and Software
All measurements are obtained as the average value over ten runs on a machine equipped with two AMD EPYC Rome 7402 CPUs (48 threads running at 2.8 GHz) and 512 GB of RAM, using Python 3.9, when not specified otherwise. The machine also features a NVIDIA A100 GPU with 40 GB of RAM and CUDA 12.1. The Ray engine was configured with default settings, resulting in 48 worker threads running. The Dask engine was configured comparably, leading to the creation of 8 workers and 48 threads for each execution. To ensure accuracy and avoid warm-up overhead, the assessment of execution time occurs once the system (e.g., JVM) has completed its warm-up process. To assess the performance of the libraries across machines with different hardware specifications, we simulated three distinct machine configurations, as detailed in Table 4. Finally, we evaluated scalability on incremental samples of Taxi and Patrol.
4. Evaluation Results
In this section, we report and analyze the results of the extensive experimental evaluation of the presented dataframe libraries. Our main goal is to provide data scientists and practitiones with useful insights for supporting them in the selection of the best solution for their data preparation tasks. Thus, our comparison is designed to assess the performance of dataframe libraries based on the operations to carry out (considering both the distinct preparators and the benefits introduced by lazy evaluation, when supported), the size and the features of the dataset at hand, and the configuration of the machine on which the pipeline is executed.
@@ 4. Evaluation Results
Figure 1. Average runtime for each stage of the data preparation pipeline (EDA, DT, DC) on different datasets.
In particular, we organize our evaluation in multiple subsections, each designed to answer one of the following research questions:
  1. Q1.
    What is the performance of the dataframe libraries in running data preparation pipelines on datasets of different size and features? (Section 4.1)
  2. Q2.
    How does lazy evaluation impact on the performance of the libraries that support it? (Section 4.2)
  3. Q3.
    How do libraries scale by varying the size of the dataset and the configuration of the underlying machine (i.e., from laptop to server)? (Section 4.3)
  4. Q4.
    How do libraries perform on the standard queries of the TPC-H benchmark? (Section 4.4)
4.1. Evaluation on Data Preparation Pipelines
{mdframed}
Summary—For EDA, Polars is generally the best performer, up to 100 times faster than Pandas. For DT, if a GPU is available, CuDF generally outperforms other libraries. For DC, ModinR achieves notable results on the largest datasets. Finally, despite not supporting the Parquet format, DataTable is the fastest library for I/O.
For each of the selected datasets (Athlete, Loan, Patrol, Taxi), Figure 1 shows the runtime for the execution of three main stages of the data preparation pipeline: EDA, DT, and DC. The I/O stage is considered separately in Figures 4 and 4. Figure 1 considers each stage in its entirety, allowing to perform lazy evaluation at stage level when supported. Additionally, in Figure 2 we present the independent performance of each distinct preparator (i.e., we force the execution separately for each preparator), normalized with respect to Pandas (the de facto standard library for managing dataframes). For each preparator, the red line represents therefore Pandas as the baseline, while distinct markers denote the speedup for each library, which is defined as follows:
$speedup=\frac{Time{\langle\textrm{Pandas},\textit{{prep}}\rangle}}{Time{\langle\textit{{lib}},\textit{{prep}}\rangle}}$
where $Time{\langle\textit{{lib}},\textit{{prep}}\rangle}$ is the time required by the library lib to run the preparator prep. Thus, a value above the red line denotes that the library outperforms Pandas; otherwise, it takes more time than Pandas to run the preparator. The logarithmic scale enables to easily assess the order of magnitude of the improvement (if any) of each library over Pandas.
@@ 4.1. Evaluation on Data Preparation Pipelines
Figure 2. Speedup over Pandas for each preparator on different datasets.
Explorative Data Analysis (EDA)
For EDA, Polars clearly stands out as the best performer on all datasets but Patrol, where CuDF outperforms other dataframe libraries (Figures 1(a)-1(d)).
In this stage, four preparators alone account for 95% of the total execution time: locate missing values (isna), locate outliers (outlier), search by pattern (srchptn), and sort values (sort). Note that Polars generally registers the best performance for two of them, locate missing values and locate outliers (Figure 2), where it is up to ${10^{4}}$ times faster than Pandas and 0.6 times faster than the second best performer, DataTable. In fact, Polars can quickly locate missing values by exploiting the specific metadata stored by Arrow arrays [^24] https://pola-rs.github.io/polars/user-guide/expressions/null/. Nevertheless, DataTable is the fastest in locating missing values on Patrol (Figure 2(c)), with a speedup of ${10^{4}}$ over Pandas and ${10^{3}}$ over Polars, thanks to its implementation of binary-search-based filtering. As highlighted by Table 3, Patrol is mostly composed of string attributes. For this type of data, Polars uses Arrow large_utf8 buffers, which are more expensive to filter than Python strings/chars, requiring therefore extra work.
For the other two most relevant preparators, in Figure 2(c) we can observe that Vaex excels in searching by pattern, being about ${10^{3}}$ times faster than Pandas and 10 times faster than the second best performer, CuDF. Vaex tends to be very efficient in column-wise operations due to its internal columnar organization of the data. On the other hand, it is much less efficient when applying a row-wise preparator, for instance in locating outliers, where it is even about 100 times slower than Pandas. Conversely, when locating missing values, Vaex shows relatively good performance on dense datasets such as Taxi and Patrol, though it is among the less efficient on sparse datasets like Loan, where its performance is does not provide improvements over Pandas (Figure 2(b)).
For sorting values, CuDF and Polars are the fastest libraries overall. SparkSQL shows remarkable advantages over Pandas as the datasets grow larger, while SparkPD maintains similar performance to the baseline. Note that in Athlete, Modin converts by default the dataframe back and forth to Pandas for the sorting operation on multiple columns, hence in this case it is the worst performer, up to 100 times slower than Pandas.
Finally, for the remaining preparators, get dataframe statistics (stats) and query columns (query), Modin outperforms other libraries on larger datasets exploiting its multithreading power, which even makes ModinD 10 times faster than Polars when running the entire stage on Taxi. Overall, Pandas maintains good performance throughout the stage.
@@
Figure 3. Average runtime for reading CSV and Parquet files.
(e) Athlete, Write
(f) Loan, Write
(g) Patrol, Write
(h) Taxi, Write
Figure 4. Average runtime for writing CSV and Parquet files.
Data Trasformation (DT)
In this stage, CuDF emerges as the best performer on Loan (Figure 1(f)) and Patrol (Figure 1(g)), while SparkSQL and Polars register the best performance (also thanks to lazy evaluation) on Athlete (Figure 1(e)) and Taxi (Figure 1(h)), respectively. The most time intensive preparators in this stage are categorical encoding (catenc), calculate new column (apply), and group dataframe (group). For catenc and group, CuDF always shows the best performance, being at least up to 100 times faster than the second best performer (Polars and ModinR, respectively). Once again, Vaex consistently outperforms the other libraries on column-wise operations (e.g., apply), being ${10^{3}}$ times faster than the second best performer CuDF (Figure 2(d)). On the other hand, this causes a drastic drop in its performance when moving to operations such as group, join, or pivot.
As in the previous stage, SparkPD results as the worst performer on several preparators. This is mainly due to the latency introduced by translating Pandas operations into Spark execution plans. Since SparkSQL does not require this translation, it often achieves significantly better performance.
When datasets scale up, libraries like ModinD and ModinR, designed for resource-intensive applications, exhibit superior performance due to their optimized handling of large-scale data compared to their performance on smaller datasets.
Data Cleaning (DC)
When moving to data cleaning operations, CuDF consistently outperforms other libraries on all datasets, except for Taxi (Figure 1(l)), where Vaex stands out as the best performer (with remarkable results achieved by Polars and Modin with both engines).
@@
The pipelines designed for the smallest datasets (Athlete and Loan) include a preparator, deduplicate rows (dedup), accounting for 70% of the overall execution time for the stage. Pandas register good results on these datasets, especially on Athlete, where only CuDF performs better (Figure 1(i)). On Loan, it is possible to notice how Vaex equals or (even significantly) exceeds the performance of CuDF for multiple preparators (Figure 2(b)), but the greater overall impact of the ones where it is outperformed by CuDF (especially the great difference in deduplicating rows) leads to a consistent difference in the overall execution time for this stage (Figure 1(j)).
@@
Results change when moving to larger datasets (Figures 1(k) and 1(l)). CuDF excels on Patrol, where it is over ${10^{3}}$ times faster than Pandas for both considered preparators, i.e., delete empty rows (dropna) and change date & time format (chdate), slightly outperforming ModinR. Nevertheless, it struggles on Taxi, where Vaex achieves the best results instead, followed closely by Modin (with both engines) and Polars.
@@
Figure 5. Average runtime for the execution of the entire data preparation pipelines.
Input/Output (I/O)
Loading data represents a fundamental step for the success of any data-driven process. In particular, the CSV format has established itself as the de facto standard for reading and writing raw data, and Pandas is recognized as an efficient and effective tool for its ingestion (Vitagliano et al., 2023). Nowadays, the diffusion of Arrow is pushing towards an increasing adoption of the Parquet format [^25] https://parquet.apache.org, known for its efficient data compression. In particular, developers often load Parquet data into memory and convert it to Arrow (Liu et al., 2023), recognized as an ideal in-memory transport layer for such data [^26] https://arrow.apache.org/docs/python/parquet.html.
Figures 4 and 4 show performance in read and write operations on both file formats, CSV and Parquet. DataTable, which maps the file into memory and navigates through it using pointers, records the best overall performance in reading CSV files, but it does not support the Parquet format. Polars is the best performer in reading Parquet files due to its capability to load them directly into memory using Arrow. When dealing with small datasets like Athlete, Parquet and CSV often exhibit almost the same performance, unlike Polars and SparkSQL (Figure 3(a)). In some cases (e.g., Vaex or CuDF), the use of Parquet can even lead to worse performance, due to the need for decompressing the data. When the size increases, the use of Parquet generally produces better results; for instance, reading Parquet files with Polars is over 100 times faster than reading CSV files (Figure 3(d)).
Write operations exhibit a wide range of performance, mainly due to the need of compressing data before saving. Parquet files come with additional metadata and encoding, which may introduce some overhead for the writing process. Nevertheless, Parquet performs better than CSV across all datasets as it integrates various efficient compression and encoding approaches, while CSV is plain text and uncompressed (Liu et al., 2023).
Table 4. Specifications of each machine configuration.
@@
4.2. Impact of Lazy Evaluation
{mdframed}
SummaryCuDF significantly outperforms other libraries when running the entire pipeline, while SparkSQL achieves remarkable results on all datasets. Lazy evaluation brings benefits mostly on large datasets, improving performance by 30% on average (up to almost 70% in some cases) compared to its eager counterpart.
Figure 5 reports for each library the runtime for the execution of the entire data preparation pipeline. For libraries supporting lazy evaluation (i.e., SparkPD, SparkSQL, and Polars), we also point out the difference compared to eager evaluation.
Considering the overall performance on the entire pipeline, CuDF consistently outperforms other libraries by effectively exploiting the power of GPU, with the only exception of the smallest dataset (Figure 5(a)). Pandas performs quite well on smaller datasets, but struggles on larger ones due to its eager approach (which requires computing intermediate results) causing higher memory consumption (Sinthong and Carey, 2021). Vaex and SparkSQL perform well on Patrol thanks to parallel processing, which allows them to efficiently exploit available CPU cores for handling many string columns.
On one hand, it can be noticed how long logical plans can limit the optimizations introduced by lazy evaluation. Yet, the execution plan overhead of SparkSQL does not yield substantial improvements for either smaller or larger datasets, with an improvement of 30% on Taxi and almost no changes in performance on Patrol. On the other hand, libraries like SparkPD leverage Pandas-specific optimizations (e.g., vectorized operations and memory efficiency) that translate into a 70% performance improvement. Further, Polars generally demonstrates good performance. In many cases, the eager API internally utilizes the lazy API before immediately collecting the result. Lazy execution in Polars leverages techniques such as streaming processing, early filtering, and projection pushdown, achieving a performance improvement of up to 15%. However, it is important to notice that these optimizations are not always available, hence in some cases lazy operations may perform comparably to their eager counterpart (Figures 5(b) and 5(d)).